# MobileCLIP: Fast Image-Text Models through Multi-Modal Reinforced Training

### v1

@@ Abstract
Contrastive pretraining of image-text foundation models, such as CLIP, demonstrated excellent zero-shot performance and improved robustness on a wide range of downstream tasks. However, these models utilize large transformer-based encoders with significant memory and latency overhead which pose challenges for deployment on mobile devices. In this work, we introduce MobileCLIP – a new family of efficient image-text models optimized for runtime performance along with a novel and efficient training approach, namely multi-modal reinforced training. The proposed training approach leverages knowledge transfer from an image captioning model and an ensemble of strong CLIP encoders to improve the accuracy of efficient models. Our approach avoids train-time compute overhead by storing the additional knowledge in a reinforced dataset. MobileCLIP sets a new state-of-the-art latency-accuracy tradeoff for zero-shot classification and retrieval tasks on several datasets. Our MobileCLIP-S2 variant is 2.3$\times$ faster while more accurate compared to previous best CLIP model based on ViT-B/16. We further demonstrate the effectiveness of our multi-modal reinforced training by training a CLIP model based on ViT-B/16 image backbone and achieving +2.9% average performance improvement on 38 evaluation benchmarks compared to the previous best. Moreover, we show that the proposed approach achieves 10$\times$-1000$\times$ improved learning efficiency when compared with non-reinforced CLIP training.
@@ 2 Related Work
One can improve learning efficiency through utilizing an enhanced training objective. Examples include image masking [70, 17, 54, 36], unimodal self-supervision [42, 34], fine-grained image-text alignment [71], contrastive learning in image-text-label space [68], and pairwise Sigmoid loss [76]. Recently, CLIPA [33] proposed training at multi-resolutions for cost-effective CLIP training. These methods are complementary to our proposed method and can be exploited for further improvements.
@@
Previous works explored extending unimodal knowledge distillation [25] to vision-language models. DIME-FM [55] proposes using in-domain unimodal data for distillation with a focus on zero-shot classification. TinyCLIP [67] trains compact CLIP models via cross-modal affinity mimicking and weight inheritance. Multi-modal distillation is also explored in setups where the student is a fused vision-language model for specific tasks [64, 30, 63]. Our proposed multi-modal reinforced training also includes cross-modal affinity mimicking [67] toward targets that are added to our reinforced datasets. Further, we extend unimodal model ensembling [32, 45] to multimodal setup, and store targets obtained from an ensemble of CLIP models.
@@ 3 Multi-Modal Reinforced Training
Our multi-modal reinforced training leverages knowledge transfer from an image captioning model and a strong ensemble of pretrained CLIP models for training the target model. It consists of two main components: i) leveraging the knowledge of an image captioning model via synthetic captions, and ii) knowledge distillation of image-text alignments from an ensemble of strong pre-trained CLIP models. We follow the dataset reinforcement strategy of [14] and store the additional knowledge (synthetic captions and teacher embeddings) in the dataset (see Fig. 3), thereby avoiding any additional training time computational overhead such as evaluating the captioning model or the ensemble teacher. The proposed training strategy results in significant improvement in learning efficiency, i.e., reaching to certain target performance with less training budget and fewer samples.
[Uncaptioned image] Figure 3: Illustration of multi-modal dataset reinforcement with one image augmentation and one synthetic caption. In practice, we use multiple image augmentations and synthetic captions.
[Uncaptioned image] Figure 4: Architecture of convolutional and reparameterizable blocks, called Text-RepMixer used in MobileCLIP’s text encoder - MCt.
@@ 3.1 Dataset Reinforcement
where ${{\bm{a}}}^{(i,j)}$ are the augmentation parameters that are sufficient to reproduce ${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}$ from ${{\bm{x}}}_{\text{img}}^{(i)}$ (see Fig. 3a). Ablations on the number and different kinds of augmentations used per image are provided in Tabs. 2(a) and 11, respectively.
@@ 3.2 Training
${\mathcal{L}_{\text{Total}}}(\mathcal{B})$ $=\lambda{\mathcal{L}_{\text{CLIP}}}(\mathcal{B})+(1-\lambda){\mathcal{L}_{\text{Distill}}}(\mathcal{B}),$ (2)
@@ 4.1 Text Encoder
CLIP [46] model paired the vision transformer with a classical transformer comprising of self-attention layers for text encoding. While this model is effective, smaller and more efficient models are preferred for mobile deployment. Recently, works like [66] have shown that convolutions can be just as effective for text encoding. In contrast, we found that purely convolutional architectures significantly underperform their transformer counterparts. Instead of using a fully convolutional architecture for text encoding, we introduce a hybrid text encoder which makes use of 1-D convolutions and self-attention layers. We design variants of MobileCLIP using both transformer and hybrid text encoders.
For hybrid text encoder, we introduce Text-RepMixer, a convolutional token mixer that decouples train-time and inference-time architectures. Text-RepMixer is inspired by reparameterizable convolutional token mixing (RepMixer) introduced in [61]. At inference, skip connections are reparameterized. The architecture is shown in Fig. 4. For Feed-Forward Network (FFN) blocks, we augment linear layers with an additional depthwise 1-D convolution of similar kernel dimensions as the token mixer, to obtain ConvFFN blocks. This structure is similar to the convolutional blocks used in [19], the main difference being the use of batchnorm and the ability to fold it with the succeeding depthwise 1-D convolutional layer for efficient inference. For all depthwise 1-D convolutions, we use a kernel size of 11. This was chosen based on accuracy-latency tradeoff, full ablation can be found in Appendix E. In order to find the optimal design for our hybrid text encoder, we start with a purely convolutional text encoder and start replacing convolutional blocks systematically with self-attention layers (see Tab. 3(b)).
@@ 4.2 Image Encoder
In FastViT, an MLP expansion ratio of 4.0 is used for FFN blocks. Recent works like [38, 67] exposed the significant amount of redundancy in linear layers of FFN block. To improve parameter efficiency, we simply lower the expansion ratio to 3.0 and increase the depth of the architecture. By doing so, we retain the same number of parameters in the image encoder. The stage configuration for the three variants are described in Tab. 3(a). MCi0 has similar stage configuration as [60]. MCi1, is a deeper version of MCi0 and MCi2 is a wider version of MCi1. The stage compute ratios in our variants are similar to [60]. We find that this design has a minimal impact on latency, but a good improvement in capacity of the model as reflected in the downstream task performance. When trained from scratch on ImageNet dataset for image classification task, MCi2 attains the same top-1 accuracy of 84.5% as FastViT [61] (previous state-of-the-art hybrid vision transformer), while being 15% smaller and 14.3% faster. More details can be found in Appendix A.
@@ 5 Experiments
In this section, we present our experimental setup, ablations on our proposed method and fast MobileCLIP architectures, and results.
@@
We have two setups for ablations and large-scale experiments. For ablations, we train on datasets with 12.8M image-text pairs using a global batch size of 8,192 and 8$\times$×NVIDIA-A100-80GB GPUs for 30-45k iterations. For large-scale training, we use a global batch size of 65,536 with 256$\times$×A100 GPUs for 200k iterations. All models are trained from scratch (see details in Appendix A).
@@
We reinforce the DataComp dataset using our multi-modal dataset reinforcement strategy. In particular, we create DataCompDR-1B and DataCompDR-12M by reinforcing DataComp-1B and DataCompDR-12M. We have a one-time generation process, the cost of which is amortized over multiple architectures and extensive ablations. We generate 5 synthetic captions per image using the coca_ViT-L-14 model in OpenCLIP [28], and strong random image augmentations (10 for DataCompDR-1B and 30 for DataCompDR-12M). We compute embeddings of an ensemble of two strong teachers (ViT-L-14 with pretrained weights datacomp_xl_s13b_b90k and openai in OpenCLIP) on augmented images as well as real and synthetic captions. Embeddings are 1536-D concatenations of 2$\times$768-D vectors. We store all reinforcements using lossless compression and BFloat16. We analyze all of our choices in Sec. 5.1.
@@ 5.1 Ablation Studies
$\lambda$
Syn. Captions
Strong Aug.
Ens. Teacher
IN-val
Flickr30k
0
44.5
41.8
0
51.9
69.3
1
54.5
66.1
1
59.3
70.5
1
61.7
72.0
0.7
60.7
74.2
@@
In contrast to uni-modal supervised and self-supervised methods for vision where strong augmentations are used [13, 59], CLIP training recipes [46] often use light image augmentations to avoid image-text mis-alignment. However, several works [45, 14, 2] demonstrated the efficacy of strong augmentations in a distillation setup. In Tab. 1 we show that strong image augmentations improve distillation performance (+4.8% on IN-val and +4.4% on Flickr30k). We provide detailed ablation on the effect of image augmentations in Appendix B.
@@
We find that using an ensemble of strong CLIP models as a teacher in our multi-modal reinforced training is crucial to achieving +2.4% IN-val improvement (Tab. 1). We also observe that the most accurate models are not the best teachers. See Appendix C for a comprehensive analysis of different teacher models.
@@
Dataset Image Text Syn.
Aug. Params
Text Emb.
Image Emb.
Size (TBs)
DataComp-12M
@@
0.9
DataCompDR-12M
5+1
30
1.9
DataComp-1B
90
DataCompDR-1B
5+1
10
140
@@
Dataset
$\mathcal{B}\in$
${\mathcal{L}_{\text{CLIP}}}$
${\mathcal{L}_{\text{Distill}}}$
Stored Syn. Caption
Stored Embeddings
Time (hours)
DataComp-12M
$\{\mathcal{B}_{\text{real}}\}$
1.3
$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
21.1
$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
4.1
DataCompDR-12M
$\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$
1.3
@@
(b) Effect of the number of self-attention layers in MCt.
Table 4: Ablation on architecture. We train for 30k iterations.
@@
We generate multiple image augmentations and synthetic captions and store them efficiently along with the teacher embeddings. We investigate the effectiveness of the number of augmentations and synthetic captions in Tabs. 2(a) and 2(b). We train models with up to 30 image augmentations and 5 synthetic captions for 45k iterations ($\sim$30 epochs). We observe that the performance nearly saturates at 5 augmentations and 2 synthetic captions suggesting each augmentation can be reused multiple times before the added knowledge is fully learned by the model. When needed, fewer augmentations and synthetic captions can help reduce the generation time and storage overhead. For maximal performance, we reinforce DataCompDR-12M and DataCompDR-1B with 10 and 30 augmentations, respectively, and 5 synthetic captions.
@@
We report the storage requirements for our reinforced datasets compared with the original DataComp dataset. In general, the storage size of datasets depends on the file format and the tradeoff between load time and the compression rate. We report the storage size of one file per image-text pair. If present, we store all corresponding reinforcements in the same file. We store files in the Pickle format and compress each file with Gzip compression. The image-text embeddings are saved in BFloat16. We report the total storage size for 12.8M samples of DataCompDR-12M and 1.28B samples of DataCompDR-1B in Tab. 2(c). We provide analysis on additional size reductions in Appendix D and verify that using BFloat16 does not impact the accuracy. For minimal storage overhead, we recommend 5 augmentations/synthetic captions for 30 epochs on DataCompDR-12M and 2 for 10 epochs on DataCompDR-1B which are based on our ablations in Tabs. 2(a) and 2(b).
@@
We ablate over the number of Text-RepMixer blocks that can effectively replace self-attention layers with negligible impact on zero-shot performance. For this ablation, we choose a 6-layer purely convolutional text encoder and systematically introduce self-attention layers in the middle. From Tab. 3(b), we find that even introducing a single self-attention layer substantially improves the zero-shot performance. The best tradeoff is with 2 blocks of Text-RepMixer and 4 blocks of self-attention layers. This variant, MCt, obtains similar performance as the pure transformer variant, while being 5% smaller and 15.8% faster.
@@ 5.2 Small Scale Regime
CLIP-B/16 [46, 42] CC-12M [4]
0.39B
11.5 + 3.3
36.5
CLIP-B/16 [46, 42] YFCC-15M [56]
0.37B
37.6
SLIP-B/16 [42] CC-12M [4]
0.39B
40.7
SLIP-B/16 [42] YFCC-15M [56]
0.37B
42.8
CLIP-B/16 DataComp-12M [18]
0.37B
10.4 + 3.3
50.1
MobileCLIP-B DataCompDR-12M
0.37B
10.4 + 3.3
65.3
CLIP-B/32 [46, 7] YFCC-15M [56]
0.49B
5.9 + 3.3
32.8
SLIP-B/32 [42, 7]
34.3
FILIP-B/32 [71, 7]
39.5
DeCLIP-B/32 [34]
43.2
DeFILIP-B/32 [7]
45.0
RILS-B/16 [70] LAION-20M [50]
0.5B
11.5 + 3.3
45.0
TinyCLIP-8M/16 [67] YFCC-15M [56]
0.75B
2.0 + 0.6
41.1
SLIP-B/16 [42] YFCC-15M [56]
0.75B
11.5 + 3.3
44.1
CLIP-B/16 DataComp-12M [18]
0.74B
10.4 + 3.3
53.5
MobileCLIP-S0 DataCompDR-12M
0.74B
1.5 + 1.6 59.1
TinyCLIP-39M/16 [67] YFCC-15M [56]
0.75B
5.2 + 1.9
63.5
MobileCLIP-S2 DataCompDR-12M
0.74B
3.6 + 3.3 64.6
MobileCLIP-B DataCompDR-12M
0.74B
10.4 + 3.3
69.1
SLIP-B/16 [42] YFCC-15M [56]
1.5B
11.5 + 3.3
45.0
CLIP-B/16 DataComp-12M [18]
1.48B
10.4 + 3.3
55.7
MobileCLIP-B DataCompDR-12M
1.48B
10.4 + 3.3
71.7
CLIPA-B/16 [33] LAION-400M [50]
${}^{\dagger}$
11.5 + 3.3
63.2
Table 5: Small-scale CLIP training. ${}^{\dagger}$ refers to multi-resolutions. Models are grouped based on the number of samples seen.

### v2

@@ Abstract
Contrastive pretraining of image-text foundation models, such as CLIP, demonstrated excellent zero-shot performance and improved robustness on a wide range of downstream tasks. However, these models utilize large transformer-based encoders with significant memory and latency overhead which pose challenges for deployment on mobile devices. In this work, we introduce MobileCLIP – a new family of efficient image-text models optimized for runtime performance along with a novel and efficient training approach, namely multi-modal reinforced training. The proposed training approach leverages knowledge transfer from an image captioning model and an ensemble of strong CLIP encoders to improve the accuracy of efficient models. Our approach avoids train-time compute overhead by storing the additional knowledge in a reinforced dataset. MobileCLIP sets a new state-of-the-art latency-accuracy tradeoff for zero-shot classification and retrieval tasks on several datasets. Our MobileCLIP-S2 variant is 2.3$\times$ faster while more accurate compared to previous best CLIP model based on ViT-B/16. We further demonstrate the effectiveness of our multi-modal reinforced training by training a CLIP model based on ViT-B/16 image backbone and achieving +2.9% average performance improvement on 38 evaluation benchmarks compared to the previous best. Moreover, we show that the proposed approach achieves 10$\times$-1000$\times$ improved learning efficiency when compared with non-reinforced CLIP training. Code and models are available at https://github.com/apple/ml-mobileclip
@@ 2 Related Work
One can improve learning efficiency through utilizing an enhanced training objective. Examples include image masking [71, 17, 55, 37], unimodal self-supervision [43, 35], fine-grained image-text alignment [72], contrastive learning in image-text-label space [69], and pairwise Sigmoid loss [77]. CLIPA [34] proposed training at multi-resolutions for cost-effective training. These methods are complementary to our proposed method.
@@
Previous works like DIME-FM [56], extends unimodal distillation [26] with a focus on zero-shot classification. TinyCLIP [68] trains compact CLIP models via cross-modal affinity mimicking and weight inheritance. Multi-modal distillation is also explored in setups where the student is a fused vision-language model for specific tasks [65, 31, 64]. Our proposed multi-modal reinforced training also includes cross-modal affinity mimicking [68]. Further, we extend unimodal model ensembling [33, 46] to multimodal setup, and store targets obtained from an ensemble of CLIP models.
@@ 3 Multi-Modal Reinforced Training
Our multi-modal reinforced training leverages knowledge transfer from an image captioning model and a strong ensemble of pretrained CLIP models for training the target model. It consists of two main components: i) leveraging the knowledge of an image captioning model via synthetic captions, and ii) knowledge distillation of image-text alignments from an ensemble of strong pre-trained CLIP models. We follow the dataset reinforcement strategy of [14] and store the additional knowledge (synthetic captions and teacher embeddings) in the dataset (see Fig. 3), thereby avoiding any additional training time computational overhead such as evaluating the captioning model or the ensemble teacher.
Figure 3: Illustration of multi-modal dataset reinforcement with one image augmentation and one synthetic caption. In practice, we use multiple image augmentations and synthetic captions.
Figure 4: Architecture of convolutional and reparameterizable blocks, called Text-RepMixer used in MobileCLIP’s text encoder MCt.
@@ 3.1 Dataset Reinforcement
where ${{\bm{a}}}^{(i,j)}$ are the augmentation parameters that are sufficient to reproduce ${\hat{{\bm{x}}}}_{\text{img}}^{(i,j)}$ from ${{\bm{x}}}_{\text{img}}^{(i)}$ (see Fig. 3a). Ablations on the number and different kinds of augmentations used per image are provided in Tabs. 4(a) and 13, respectively.
@@ 3.2 Training
${\mathcal{L}_{\text{Total}}}(\mathcal{B})$ $=(1-\lambda){\mathcal{L}_{\text{CLIP}}}(\mathcal{B})+\lambda{\mathcal{L}_{\text{Distill}}}(\mathcal{B}),$ (2)
@@ 4.1 Text Encoder
CLIP [47] model paired the vision transformer with a classical transformer comprising of self-attention layers for text encoding. While this model is effective, smaller and more efficient models are preferred for mobile deployment. Recently, works like [67] have shown that convolutions can be just as effective for text encoding. In contrast, we found that purely convolutional architectures significantly underperform their transformer counterparts. Instead of using a fully convolutional architecture for text encoding, we introduce a hybrid text encoder which makes use of 1-D convolutions and self-attention layers.
For hybrid text encoder, we introduce Text-RepMixer, a convolutional token mixer that decouples train-time and inference-time architectures. Text-RepMixer is inspired by reparameterizable convolutional token mixing (RepMixer) introduced in [62]. At inference, skip connections are reparameterized. The architecture is shown in Fig. 4. For Feed-Forward Network (FFN) blocks, we augment linear layers with an additional depthwise 1-D convolution of similar kernel dimensions as the token mixer, to obtain ConvFFN blocks. This structure is similar to the convolutional blocks used in [20], the main difference being the use of batchnorm and the ability to fold it with the succeeding depthwise 1-D convolutional layer for efficient inference. The design choices for Text-RepMixer is discussed in Appendix F. In order to find the optimal design for our hybrid text encoder, we start with a purely convolutional text encoder and start replacing convolutional blocks systematically with self-attention layers (see Tab. 5). Tab. 1, show the efficacy of our text encoder when compared with CLIP’s base text encoder. Our model is smaller, faster and obtains similar performance as the larger base text encoder when paired with efficient backbones like ViT-S/16.
@@ 4.2 Image Encoder
In FastViT, an MLP expansion ratio of 4.0 is used for FFN blocks. Recent works like [39, 68] exposed the significant amount of redundancy in linear layers of FFN block. To improve parameter efficiency, we simply lower the expansion ratio to 3.0 and increase the depth of the architecture. By doing so, we retain the same number of parameters in the image encoder. The stage configuration for the three variants are described in Appendix A. MCi0 has similar stage configuration as [61]. MCi1, is a deeper version of MCi0 and MCi2 is a wider version of MCi1. The stage compute ratios in our variants are similar to [61]. We find that this design has a minimal impact on latency, but a good improvement in capacity of the model, reflected in the downstream task performance, see Appendix B. In Tab. 1, we compare our MCi encoder with a similar sized FastViT-MA36 when used as image encoders in a CLIP model. Our model obtains much better zero-shot IN-val performance while being 16.3% faster.
Text Enc. Latency (txt) 0-shot IN-val
Base 3.3 53.4
MCt (Ours) 1.6 53.6
Image Enc. Latency (img) 0-shot IN-val
FastViT-MA36 4.3 58.9
MCi2 (Ours) 3.6 60.0
Table 1: (a) Base vs. MCt text encoders with ViT-S/16. (b) FastViT vs. MCi image encoders with Base text encoder. Trained for 30k iters ($\sim$0.24B seen samples) on DataCompDR-12M.
@@ 5 Experiments
In this section, we present our experimental setup and results.
@@
We have two setups for ablations and large-scale experiments. For ablations, we train on datasets with 12.8M image-text pairs using a global batch size of 8,192 and 8$\times$×NVIDIA-A100-80GB GPUs for 30-45k iterations. For large-scale training, we use a global batch size of 65,536 with 256$\times$×A100 GPUs for 200k iterations. All models are trained from scratch (see details in Appendix B).
@@
We reinforce the DataComp dataset using our multi-modal dataset reinforcement strategy. In particular, we create DataCompDR-1B and DataCompDR-12M by reinforcing DataComp-1B and DataCompDR-12M. We have a one-time generation process, the cost of which is amortized over multiple architectures and extensive ablations. We generate 5 synthetic captions per image using the coca_ViT-L-14 model in OpenCLIP [29], and strong random image augmentations (10 for DataCompDR-1B and 30 for DataCompDR-12M). We compute embeddings of an ensemble of two strong teachers (ViT-L-14 with pretrained weights datacomp_xl_s13b_b90k and openai in OpenCLIP) on augmented images as well as real and synthetic captions. Embeddings are 1536-D concatenations of 2$\times$768-D vectors. We store all reinforcements using lossless compression and BFloat16. We analyze all of our choices in Sec. 5.1. One seen sample for DataCompDR is a triplet of one randomly augmented image, one ground-truth caption, and one randomly picked synthetic caption.
@@ 5.1 Ablation Studies
$\lambda$ Syn. Captions Strong Aug. Ens. Teacher IN-val Flickr30k
0 44.5 41.8
0 51.9 69.3
1 54.5 66.1
1 59.3 70.5
1 61.7 72.0
0.7 60.7 74.2
@@
In contrast to uni-modal supervised and self-supervised methods for vision with strong augmentations [13, 60], CLIP training recipes [47] often use light image augmentations to avoid image-text mis-alignment. However, several works [46, 14, 2] demonstrated the efficacy of strong augmentations in a distillation setup. In Tab. 2 we show that strong image augmentations improve distillation performance (+4.8% on IN-val and +4.4% on Flickr30k). We provide detailed ablation on the effect of image augmentations in Appendix C.
@@
We find that using an ensemble of strong CLIP models as a teacher in our multi-modal reinforced training is crucial to achieving +2.4% IN-val improvement (Tab. 2). We also observe that the most accurate models are not the best teachers. See Appendix D for a comprehensive analysis of different teacher models.
@@
Dataset Image Text Syn. Aug. Params Text Emb. Image Emb. Size (TBs)
DataComp-12M 0.9
@@
DataCompDR-12M 5+1 30 1.9
DataComp-1B 90
DataCompDR-1B 5+1 10 140
@@
Dataset $\mathcal{B}\in$ ${\mathcal{L}_{\text{CLIP}}}$ ${\mathcal{L}_{\text{Distill}}}$ Stored Syn. Caption Stored Embeddings Time (hours)
DataComp-12M $\{\mathcal{B}_{\text{real}}\}$ 1.3
- $\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$ 21.1
- $\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$ 4.1
DataCompDR-12M $\{\mathcal{B}_{\text{real}}$, $\mathcal{B}_{\text{syn}}\}$ 1.3
@@
Table 5: Ablation on architecture. Effect of the number of self-attention layers in MCt. We train for 30k iterations.
@@
We generate multiple image augmentations and synthetic captions and store them efficiently along with the teacher embeddings. We investigate the effectiveness of the number of augmentations and synthetic captions in Tabs. 4(a) and 4(b). We train models with up to 30 image augmentations and 5 synthetic captions for 45k iterations ($\sim$30 epochs). We observe that the performance nearly saturates at 5 augmentations and 2 synthetic captions suggesting each augmentation can be reused multiple times before the added knowledge is fully learned by the model. When needed, fewer augmentations and synthetic captions can help reduce the generation time and storage overhead. For maximal performance, we reinforce DataCompDR-12M and DataCompDR-1B with 10 and 30 augmentations, respectively, and 5 synthetic captions.
@@
We report the storage requirements for our reinforced datasets compared with the original DataComp dataset. We report the storage size of one file per image-text pair. If present, we store all corresponding reinforcements in the same file. We store files in the Pickle format and compress each file with Gzip compression. The image-text embeddings are saved in BFloat16. We report the total storage size for 12.8M samples of DataCompDR-12M and 1.28B samples of DataCompDR-1B in Tab. 4(c). We provide analysis on additional size reductions in Appendix E and verify that using BFloat16 does not impact the accuracy. For minimal storage overhead, we recommend 5 augmentations/synthetic captions for 30 epochs on DataCompDR-12M and 2 for 10 epochs on DataCompDR-1B which are based on our ablations in Tabs. 4(a) and 4(b).
@@
We ablate over the number of Text-RepMixer blocks that can effectively replace self-attention layers with negligible impact on zero-shot performance. For this ablation, we choose a 6-layer purely convolutional text encoder and systematically introduce self-attention layers in the middle. From Tab. 5, we find that even introducing a single self-attention layer substantially improves the zero-shot performance. The best tradeoff is with 2 blocks of Text-RepMixer and 4 blocks of self-attention layers. This variant, MCt, obtains similar performance as the pure transformer variant, while being 5% smaller and 15.8% faster.
@@ 5.2 Small Scale Regime
CLIP-B/16 [47, 43] CC-12M [4] 0.39B 11.5 + 3.3 36.5
CLIP-B/16 [47, 43] YFCC-15M [57] 0.37B 37.6
MobileCLIP-B CC-12M [4] 0.37B 10.4 + 3.3 38.1
SLIP-B/16 [43] CC-12M [4] 0.39B 11.5 + 3.3 40.7
SLIP-B/16 [43] YFCC-15M [57] 0.37B 42.8
MobileCLIP-B DataComp-12M [18] 0.37B 10.4 + 3.3 50.1
MobileCLIP-B DataCompDR-12M 0.37B 10.4 + 3.3 65.3
CLIP-B/32 [47, 7] YFCC-15M [57] 0.49B 5.9 + 3.3 32.8
SLIP-B/32 [43, 7] 34.3
FILIP-B/32 [72, 7] 39.5
DeCLIP-B/32 [35] 43.2
DeFILIP-B/32 [7] 45.0
RILS-B/16 [71] LAION-20M [51] 0.5B 11.5 + 3.3 45.0
TinyCLIP-8M/16 [68] YFCC-15M [57] 0.75B 2.0 + 0.6 41.1
SLIP-B/16 [43] YFCC-15M [57] 0.75B 11.5 + 3.3 44.1
CLIP-B/16 DataComp-12M [18] 0.74B 10.4 + 3.3 53.5
MobileCLIP-S0 DataCompDR-12M 0.74B 1.5 + 1.6 59.1
TinyCLIP-39M/16 [68] YFCC-15M [57] 0.75B 5.2 + 1.9 63.5
MobileCLIP-S2 DataCompDR-12M 0.74B 3.6 + 3.3 64.6
MobileCLIP-B DataCompDR-12M 0.74B 10.4 + 3.3 69.1
SLIP-B/16 [43] YFCC-15M [57] 1.5B 11.5 + 3.3 45.0
CLIP-B/16 DataComp-12M [18] 1.48B 10.4 + 3.3 55.7
MobileCLIP-B DataCompDR-12M 1.48B 10.4 + 3.3 71.7
CLIPA-B/16 [34] LAION-400M [51] 2.69B 11.5 + 3.3 63.2
Table 6: Small-scale CLIP training. MobileCLIP-B notation refers to our re-implementation of ViT-B/16 image encoder and standard Base text encoder. refers to multi-resolutions. Models are grouped based on the number of samples seen.